Jo's age plus mo's age is 27 and jo's age plus flo's age is 38 and mo's age plus flo's age is 33. How old are jo , mo and Flo?
Answer by ankor@dixie-net.com(22740)

write an equation for each statement
:
jo's age plus mo's age is 27
j + m = 27
:
and jo's age plus flo's age is 38
j + f = 38
:
and mo's age plus flo's age is 33.
m + f = 33
:
Add the three equations
j + m + 0 = 27
j + 0 + f = 38
0 + m + f = 33
--------------------
2j + 2m + 2f = 98
simplify, divide by 2
j + m + f = 49
We know j+m = 27
27 + f = 49
f = 49 - 27
f = 22 is flo's age
then replace f
m + 22 = 33
m = 33-22
m = 11 is Moe's age
and replace m
j + 11 = 27
j = 27 - 11
j = 16 yrs is joe's age
:
You can confirm these solutions yourself
How old are jo , mo and Flo?